5. For a standard normal curve, find the area to the left of z = -2.03. ____________

6A. For a standard normal curve, find the area to the right of z = -2.14. ____________

7B. For a standard normal curve, find the area between z = 1.21 and z = 2.31. ____________

8C. For a standard normal curve, find the area between z = -2.46 and z = 1.54.
Answer #1

solution

(A)P(z < -2.03)

Using z table  

=0.0212

area =0.0212

(B)

P(z > -2.14 ) =1 - P(z < -2.14 )

Using z table,
= 1 -0.0162

= 0.9838

area =0.9838

(C)

P( 1.21< Z <2.31 )

= P(Z < 2.31) - P(Z <1.21 )

Using z table,  

= 0.9896 - 0.8869

probability=0.1027

(D)

P( -2.46< Z <1.54 )

= P(Z < 1.54) - P(Z <-2.46 )

Using z table,  

= 0.9382 - 0.0069

probability=0.9313